Another published set of refractive indices for the eye components has ncornea = 1.376, naqueous

= 1.336, nlens = 1.40, nvitreous = 1.337, and the four radii of curvature as (from anterior to posterior) 

7.8, 6.4, 10.1, and 6.1 nm. Calculate what are the focusing power and focal length of this eye when 

the eye is in contact with water.


1
Expert's answer
2021-06-03T09:16:17-0400

"D_{cor}=D_1+D_2=\frac{n_2-n_1}{R_1}+\frac{n_3-n_2}{R_2}=\frac{1.376-1.336}{7.8}+\frac{1.337-1.336}{6.4}=5.28\cdot 10^6 ~dpt,"

"D_{cri}=D_3+D_4=\frac{n_3-n_2}{R_3}+\frac{n_4-n_3}{R_4}=\frac{1.337-1.336}{10.1}+\frac{1.400-1.337}{6.1}=108.79\cdot 10^6 ~dpt,"

"D=D_{cor}+D_{cri}=114.07\cdot 10^6 ~dpt,"

"F=\frac 1D=8.77~nm."